30. The B B and H C C and H D. D and H . The profit earned by the profie-maximiring monopolist is repesented by the A. ABGH BDEG C. BCFG D. CDEF B. 32. The profit per unit is: AF-G B. E-F CE-G D. G-H 33. The monopolists supply curve is the marginal-cost curve above average variable cost A. Tnue B. False 34. Government regulations to force a natural monopoly to charge a price equal to its marginal cost wil: A. improve profit for the natural monopolist B. raise the production of the good C. attract additional firms to enter the market D cause the monopolist to exit the market because its losing money 35. The purpose of antitrust laws is to A. regulate the prices charged by a monopoly B. increase competition in an industry by preventing mergers and breaking up large firms C. increase merger activity to help generate synergies that reduce costs and raise efficiency D. create public ownership of natural monopolies. 36. Which of the following statements about price discrimination is NOT true? A. Price discrimination can raise economic welfare. B. Perfect price discrimination transforms consumer surplus and deadweight loss to monopolists profit C. Perfect price discrimination generates a deadweight loss. D. Price discrimination increases a monopolists profit. 37. A monopoly is able to continue to generate economic profits in the long-run because: A. potential competitors sometimes dont notice the profits. B, there are barriers to entry to that market C. the monopolist is financially powerful. D. antitrust laws eliminate competitors for a specified number of years. 23
Answer #1

30. Monopolist produces according to the rule MR = MC and selects a price determined by this quantity. Hence it produces H and charges B. Select Option B

31. Profit = (P - ATC)*Q = area of rectangle BDEG. Select Option B

32. Per unit profit = P - ATC = E - G. Select Option C

33. False because a monopoly cannot have a supply curve

34. It will reduce the price and increase the quantity so option B is correct

35. B

36. C

37. B